#2. The Babaloo Corporation sells lollipops. Selling price $7.00/unit Variable expense$4.00/unit Current Sales 12,000 units Babaloo found that the current net operating income is equal to 25% of current sales. What is Babaloos fixed cost? a. $17,000 b. $21,000 .$15,000 d. $36,000 e.$25,200
Answer #1

Answer . C option - $15,000

Total Sales = 7*12000 = $84,000

Total Variable Costs = 4*12,000 = $48,000

Net Operating Income = 25% of net current sales = 0.25*84,000 =$21,000

Also we can arrive a the operating income from sales as in given formula = Sales - fixed costs - vaiable costs

Rearranging the formula i can say that fixed cost equals to = 21,000 -84000 +48000 =$15,000